1. Suppose that a Consumption equation is of the form:

C = 145 + 0.67 Yd, where Yd is disposable income.

a. Interpret what the figures 145 and 0.67 imply in the Consumption equation.
b. What is the associated Savings function? What do the intercept and slope mean?
c. Sketch the savings and the associated consumption functions.
d. Compute the level of consumption if income (Yd) is GHS 200.
1
Expert's answer
2021-10-20T10:02:29-0400

Solution:

a.). 145 – This refers to autonomous consumption, which is the expenditure that consumers must make when they have no disposable income.

0.67 – This refers to the marginal propensity to consume (MPC). It is the proportion of an aggregate raise in pay that a consumer spends on the consumption of goods and services.

 

b.). Saving function = Y – C

S = -C + (1 – b) Y

MPS = 1 – 0.67 = 0.33

Saving function = -145 + 0.33Y

The intercept refers to the autonomous consumption and the slope refers to the marginal propensity to consume and indicated induced consumption.

 

c.). The savings and the associated consumption graph are as follows:

 



d.). C = 145 + 0.67Yd

Yd = 200

C = 145 + 0.67 (200)

C = 145 + 134 = 279

The level of consumption = GHS 279